OpenStudy (anonymous):

\[\frac{ \sqrt{90} }{ \sqrt{10} }\] What is the decimal answer and the fraction answer?

OpenStudy (anonymous):

\[\frac{ \sqrt{90} }{ \sqrt{10} }\]\[\frac{ \sqrt{90}*\sqrt{10} }{ \sqrt{10}*\sqrt{10}}\]\[\frac{ \sqrt{90*10} }{ 10 }\]\[\frac{ \sqrt{900} }{ 10 }\]\[\frac{ 30 }{ 10 }\]Fraction\[\frac{ 3 }{ 1 }\] Decimal is 3.0

OpenStudy (anonymous):

Thank you!